OpenStudy (anonymous):

The sides of a triangle are in the ratio of 1:2:2 and the perimeter of the triangle is 15 feet. The length of the shortest side is 2.5 feet 3 feet 5 feet 7.5 feet

OpenStudy (ghazi):

x+2x+2x=15

OpenStudy (ghazi):

solve for x

OpenStudy (ajprincess):

Let the x be the length of the shortest side. Then the length of other two sides will be 2x each since the ratio is 1:2:2. Perimeter=sum of the lengths of the sides x+2x+2x=15

OpenStudy (ghazi):

sum of the sides give you perimeter let's say length of one side is x now sum of all the sides = x+2x+2x which is equal to 15 x+2x+2x=15 now solve for x

OpenStudy (anonymous):

do i divide first

OpenStudy (anonymous):

15/2x?

OpenStudy (ghazi):

5x=15 divide by 5 on both the sides x=3

OpenStudy (anonymous):

oh because of like terms

OpenStudy (ghazi):

\[x+2x+2x=15....5x=15.....\frac{ 5x }{ 5 }=\frac{ 15 }{ 5 }....x=3\]

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

awsome

OpenStudy (anonymous):

thanks

OpenStudy (ghazi):

:)